OpenStudy (anonymous):

the set of zeros of f(x)=^3=4x^2+4x is

OpenStudy (sloppycanada):

I assume you mean - \[f(x)=x^3=4x^2+4x?\]

OpenStudy (anonymous):

yes @sloppycanada

OpenStudy (zzr0ck3r):

that makes no sense... Do you mean \[f(x) = x^3+4x^2+4x\]?

OpenStudy (sloppycanada):

I'm really hoping it is @zzr0ck3r

OpenStudy (zzr0ck3r):

:)

OpenStudy (sloppycanada):

If it's \[x^3+4x^2+4x\], I can do that. @cutegirl

OpenStudy (anonymous):

it is @sloppycanada

OpenStudy (zzr0ck3r):

factor a \(x\) out and tell me what you got

OpenStudy (anonymous):

^3+4^2+4=8^5?

OpenStudy (anonymous):

@zzr0ck3r

OpenStudy (zzr0ck3r):

what does ^3 mean?

OpenStudy (zzr0ck3r):

where did the x's go?

OpenStudy (danjs):

factor the common x term from each f(x) = (x)*(x^2+4x+4) The zeros are found when it crosses the x axis, or when y = f(x) = 0, solve for the x values

OpenStudy (anonymous):

so 0,2 ? @DanJS

OpenStudy (anonymous):

its really 0,-2